As sales manager, Joe Batista was given the following static budget report for selling expenses in the Clothing Department of Soria Company for the month of October.

Sales in units 7500 9000 1500 Favorable
Variable expenses
Sales commissions 1500 2160 660 Unfavorable
Advertising expense 750 630 120 Favorable
Travel expense 3000 3150 150 Unfavorable
Free samples given out 1125 990 135 Favorable
Total variable 6375 6930 555 Unfavorable
Fixed expenses
Rent 1000 1000 0 N
Sales salaries 1000 1000 0 N
Office salaries 600 600 0 N
Depreciation-autos (sales staff) 400 400 0 N
Total fixed 3000 3000 0 N
Total expenses 9375 9930 555 Unfavorable

As a result of this budget report, Joe was called into the president’s office and congratulated on his fine sales performance. He was reprimanded, however, for allowing his costs to get out of control. Joe knew something was wrong with the performance report that he had been given. However, he was not sure what to do, and comes to you for advice.

Prepare a budget report based on flexible budget data to help Joe. (List variable costs before fixed costs. Do not leave any answer field blank. Enter 0 for amounts.)
Answer #1
Budget actual
Sales in units 9000 9000
Variable expenses
Sales commissions (1500/7500=0.20) 1800 2160 360 U
Advertising expense (750/7500=0.10) 900 630 270 F
Travel expense (3000/7500=0.40) 3600 3150 450 F
Free samples (1125/7500=0.15) 1350 990 360 F
Total variable expenses 7650 6930 720 F
Fixed expenses
Rent 1000 1000 0
Sales Salaries 1000 1000 0
Office salaries 600 600 0
Depreciation 400 400 0
Total fixed expenses 3000 3000 0
Total expenses 10650 9930 720 F